6. An object that sinks in water has a mass in air of 0.0675 kg. Its apparent mass when...

An object that sinks in water has a mass in air of 0.0675 kg. Its apparent mass when submerged in water is 0.0424 kg. What is the specific gravity SG of the object?
Answer #1
The mass of water with the volume of the object is 0.0675 - 0.0424 = 0.0251kg.
The volume is V = M_w / p_w = 0.0252kg / 1e3kg/m^3 = 0.0252e-3 m^3.
So the object's density is p = m/V = 0.0675kg / 0.0252e-3m^3 = 2679 kg/m^3
Answer #2

Spo true - sample PH20

where Psample is the density of the sample and H20 is the density of water.

Since the density of water is 1 g/cm^3, the sp. g. = 0.0675/(0.0675-0.0424)= 2.689
Answer #3

The mass of water with the volume of the object is 0.0675 - 0.0424 = 0.0251kg.
The volume is V = M_w / p_w = 0.0252kg / 1e3kg/m^3 = 0.0252e-3 m^3.
So the object's density is p = m/V = 0.0675kg / 0.0252e-3m^3 = 2679 kg/m^3.